OpenStudy (anonymous):

Find the indicated outputs for f(x) = 5x^2 - 5x f(0) = ? f(-1) =? f(2) = ?

OpenStudy (anonymous):

replace x by the various numbers, see what you get. for example \[f(2)=5\times 2^2-5\times 2=5\times 4-10=20-10=10\]

OpenStudy (anonymous):

for x = -1 use parentheses

OpenStudy (aroub):

f(x) = 5x^2 -5x f(0)=5(0)-5(0)= f(-1)= 5 -1^2 -5(-1)= f(2)=5 (2)^2-5(2) = just solve now!! :D

OpenStudy (anonymous):

So f(2) =? would be 10?

OpenStudy (anonymous):

Yeap!

OpenStudy (anonymous):

So f(-1) would be 5?

OpenStudy (anonymous):

so f(0) would be -1?

OpenStudy (aroub):

no for f(0) =0

OpenStudy (aroub):

bec. 5 times zero = zero

OpenStudy (anonymous):

ah...okay I get it. thanks